Let f(x) = x^2 + 6x. Use the definition (ε − δ method) to show that
lim
x→−3 f(x) = −9
1
Expert's answer
2020-12-07T16:44:14-0500

We remind, that by definition "\lim_{x\rightarrow a}f(x)=B" means:

"\forall\epsilon>0" "\exists\delta>0": "|x-a|<\delta" "\Rightarrow" "|f(x)-B|<\epsilon" .

We have to show that "\lim_{x\rightarrow -3}f(x)=-9" . Let us fix an arbitrary "\epsilon>0" . We need to find such "\delta>0" that inequality "|x+3|<\delta" will imply "|x^2+6x+9|<\epsilon". We can rewrite the latter as: "x^2+6x+9=(x+3)^2" From the latter it is clear that we can take "\delta=\sqrt{\epsilon}" . Thus, "\forall\epsilon>0" "\exists\delta=\sqrt{\epsilon}>0" such that "|x+3|<\delta=\sqrt{\epsilon}" implies "(x+3)^2<\epsilon" .